### v1

@@ 3 Experiment Results and Main Findings
Another key finding is that the simplest degree-based greedy (Deg-Greedy) serves as a remarkably strong baseline. As shown in Tables 1, 2 and 3, leveraging neural networks for node selection, LTFT often perform comparably to Deg-Greedy, particularly on larger or denser graphs. For example on ER graphs when $n=1000$ or $3000$, LTFT gives performance within 2% of Deg-Greedy (Table 1). An exception is for BA graphs where the graph is very dense ($n=3000,m=500$) and clear patterns emerge in the graph, LTFT surpasses Deg-Greedy. Additionally, DIFUSCO and PCQO fail to outperform Deg-Greedy on larger or less sparse graphs, such as $n=3000$, $d>30$.
4 AI-Inspired vs Deg-Greedy
Our results show that the state-of-the-art AI-inspired algorithms for MIS still do not outperform the best heuristic ReduMIS. The surprising finding was that they also often do not outperform the simplest classical heuristic, Deg-Greedy, especially on large and dense graphs. In this section, we delve deeper into this comparison (Sections 4.1 and 4.2). Furthermore, we explore the impact of augmenting various algorithms with a local search as a post hoc step to enhance solution quality (Section 4.3).
@@ 4.1 Comparison between Deg-Greedy and LTFT
Figure 1: Percentage of rounds when LTFT selects the node with smallest possible degree, i.e., behaves similarly to degree-based greedy. On ER graphs when the graph is dense (closer to the bottom left corner), the percentage of rounds LTFT selects the node with smallest possible degree is higher, i.e., behaves more similarly to degree-based greedy. On BA graphs, the percentage to choose the smallest possible node is generally lower, but on denser BA graphs LTFT also behaves more like degree-based greedy.
Deg-Greedy sequentially picks nodes for the independent set. At each step, it picks the node with the smallest degree in the residual graph (where the nodes in the independent set and their neighbors are removed). It does not reverse any decisions (ie once picked, the node stays in the independent set). As in Section 2.2 we call it a non-backtracking algorithm. LTFT is also a non-backtracking algorithm and it often perform similarly to Deg-Greedy in Tables 1 and 2. It uses a trained policy network GFlowNets (Bengio et al., 2021a) to pick a node for the independent set at each step. Thus, we can naturally compare LTFT with Deg-Greedy by investigating how close this trained policy compares to the naive policy in Deg-Greedy.
The results are shown in Figure 1. Overall, we observe that LTFT frequently selects nodes with very small degrees. On ER graphs where the average degree is at least 30, LTFT picks the node with the smallest degree over 85% of rounds except for sparse graphs ($d=10$ and $(n,d)=(3000,30)$). On BA graphs, while the percentage is lower, it still exceeds 75% on large and sufficiently dense graphs. Moreover, in cases where LTFT selects the smallest degree nodes less frequently, its performance is worse than Deg-Greedy.
@@ 4.2 Serialization: allows comparing to Deg-Greedy
Figure 2: The percentage to choose the smallest possible degree node on different part of the (degree-based) serialization. We find that the best algorithms (OnlineMIS, ReduMIS, iSCO) and the best performing learning-based algorithm LwD share a similar characteristic pattern, that they have high consistency with degree-based greedy on second and third part of the serialization, while on the first part there is a low consistency. On the other hand, PCQO and DIFUSCO has low consistency with degree-based greedy in general.
@@
To evaluate the algorithms, we compute the percentage of rounds in which the smallest degree node is selected during serialization, similar to our comparison between LTFT and Deg-Greedy. Due to random tie-breaking in Algorithm 1, we repeat the process 100 times and select the serialization with the highest number of rounds selecting the smallest degree node. Although Deg-Greedy theoretically achieves 100% smallest-degree selections in its best serialization, random tie-breaking prevents us from perfectly recovering this with 100 repetitions. Instead of reporting the overall percentage, we divide the serialization into three equal parts and report the percentages for each part.
The results, shown in Figure 2, compare the percentage of smallest-degree node selections across different algorithms. Due to space constraints, we present results only for ER and BA graphs under selected parameters; additional results are available in Section C.2.
Algorithms that often perform the best, namely OnlineMIS, ReduMIS, and iSCO, exhibit a consistent pattern after serialization. In the first one-third of the serialization, these algorithms deviate significantly from selecting the smallest-degree nodes. However, in the middle and final thirds, the percentage of smallest-degree node selections increases substantially. This suggests that while degree-based greedy heuristics may appear shortsighted initially, they are highly effective in the later stages of solution construction. Interestingly, LwD, which performs the best among learning-based methods tested in our setting, also shares this pattern.
As for PCQO and DIFUSCO, they show consistently low percentages of smallest-degree node selections throughout the serialization, particularly on ER graphs.
Through serialization, we observe distinct differences in node selection patterns among algorithms. Our findings suggest that AI-based methods might fail to utilize (PCQO and DIFUSCO) or emphasize too much (LTFT) on simple yet highly effective heuristics, such as greedily selecting the smallest-degree node, which may partly explain their performance limitations.
In Section C.3, we perform a pseudo-natural serialization for LwD. LwD is also a non-backtracking algorithm like LTFT, but it does not have a “natural serialization” like LTFT (Section 4.1) because it chooses several nodes in a step. The pseudo-natural serialization performs serialization in each step of LwD. The results in Figure 6 align with our “counterfactual” serialization results here.
@@ 4.3 Incorporating local search to improve solution
In the previous sections, we show that solutions generated by AI-based algorithms generally differ from those produced by degree-based greedy methods, which may explain their inferior performance on MIS problems. A natural idea is to enhance these solutions with simple heuristics, such as local search. Local search post-processing has also been used for AI-algorithms in previous works (Ahn et al., 2020; Böther et al., 2022).
@@
We applied the $2$-improvement local search (Andrade et al., 2012) (details in B.7), which is used in KaMIS, as a post-processing step to all algorithms (except OnlineMIS and ReduMIS since they already has local search), and the resulting performance improvements are presented in Table 4.
Table 4: Incoperating local search as a post-processing procedure We add $2$-improvement local search for the algorithm tested. We report the performance after the local search and the improvement from local search in (). AI-inspired methods still don’t outperform ReduMIS after local search. We also observe that if the solution after serialization exhibits a high percentage of smallest-degree node selections in the later stages, there is relatively little room for improvement through local search. Conversely, if the solution after serialization shows a low percentage of smallest-degree node selections, there is greater potential for improvement via local search.
@@
Heuristics GPU-acc Learning-based param Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO ER 1000,30 152.75 (1.75) 158.88 163.75 163.50 (0) 148.50 (5.50) 158.62 (0.24) 151.38 (1.38) 150.62 (6.87) 1000,100 60.75 (0.13) 64.75 66.63 66.50 (0) 59.00 (3.38) 64.12 (0.24) 61.88 (1.00) 57.88 (2.50) 3000,30 456.12 (4.24) 480.88 493.13 491.62 (0) 453.88 (11.50) 474.00 (0.75) 454.38 (5.38) 442.75 (29.37) BA 1000,15 234.88 (1.63) 245.00 246.38 246.25 (0) 235.50 (4.50) 243.12 (0) 230.50 (6.25) 237.88 (1.50) 1000,50 108.38 (1.00) 115.75 116.88 116.75 (0) 110.25 (2.37) 113.12 (0.12) 106.75 (0.37) 108.75 (3.50) 3000,15 714.62 (7.50) 749.63 754.50 752.00 (0) 723.50 (12.88) 730.50 (2.75) 693.00 (31.12) 731.75 (5.37)
@@
As shown in Table 4, algorithms like PCQO and DIFUSCO benefit significantly more from the local search post-processing compared to others, such as Deg-Greedy and LwD. This observation aligns with our earlier findings: If the solution after serialization exhibits a high percentage of smallest-degree node selections in the later stages, there is relatively little room for improvement through local search. Conversely, if the solution after serialization shows a low percentage of smallest-degree node selections, there is greater potential for improvement via local search.
In addition, although all algorithms except iSCO have improvements after local search, they still perform worse than ReduMIS in most cases.
In summary, our analysis highlights a promising direction for designing machine learning-based combinatorial optimization algorithms. Rather than relying solely on end-to-end methods like PCQO or DIFUSCO, incorporating classical heuristics, such as greedily selecting the smallest-degree node, into the overall algorithm may yield better results. One potential approach could involve using machine learning algorithms to identify a small subset of nodes, followed by a degree-based greedy method to complete the solution.
4.4 KaMIS refutes famous conjecture?
For theoretical analysis for MIS on ER graphs and regular graphs, see Coja-Oghlan & Efthymiou (2015); Wormald et al. (1999); Barbier et al. (2013); Gamarnik & Sudan (2014). In ER graphs with $n$ nodes and average degree $d$, the MIS has size $\frac{2n\ln d}{d}$ for asymptotically large $n$ and $d$, and simplest random greedy achieves half-optimal at $\frac{n\ln d}{d}$ (Grimmett & McDiarmid, 1975). Yet, there is no known polynomial-time algorithm which can achieve MIS size $(1+{\epsilon})\frac{n\ln d}{d}$ for any constant ${\epsilon}$ and it is conjectured that polynomial-time algorithms cannot do better than $(1+o(1))\frac{n\ln d}{d}$ (Coja-Oghlan & Efthymiou, 2015). Thus it is natural to measure the goodness of an algorithm by the ratio of the MIS size obtained to $\frac{n\ln d}{d}$. This ratio can faciliate comparison across different $n$ and $d$’s.
Figure 3 plots this ratio for several algorithms. Surprisingly for ReduMIS and OnlineMIS this ratio is consistently larger than $1.2$ and often larger than $1.3$ even for fairly large $n,d$. These results seem to bring the conjecture in doubt, but not disprove per se since the conjecture concerns asymptotically large $n$ and $d$. But our results could encourage further analysis and potential collaboration between researchers on theoretical and empirical aspects of MIS on random graphs.
Figure 3: Heatmap for ratios of MIS size to $\frac{n\ln d}{d}$ on ER graphs. We find that the ratio for ReduMIS and KaMIS is consistently larger than 1.2, suggesting that ReduMIS and KaMIS might surpass the conjectured upper bound (Coja-Oghlan & Efthymiou, 2015).
5 Related Works
Classical and heuristic methods for MIS
Classical methods for MIS range from simple greedy algorithms to advanced solvers like KaMIS which involves a number of heuristics. There are various existing heuristics for MIS, such as reduction techniques (Butenko et al., 2002; Xiao & Nagamochi, 2013; Akiba & Iwata, 2016), local search (Andrade et al., 2012), and evolutionary algorithms (Back & Khuri, 1994; Borisovsky & Zavolovskaya, 2003; Lamm et al., 2015). KaMIS (Lamm et al., 2017; Dahlum et al., 2016; Hespe et al., 2019) was developed based on these heuristics. In addition, MIS can be formulated into a binary integer programming problem (Nemhauser & Trotter Jr, 1975), which can be solved by the state-of-art integer programming solver Gurobi Gurobi Optimization, LLC (2024).
Machine learning for combinatorial optimization
In recent years, various ML-based algorithms have been developed for the MIS problem and most of them are based on graph neural networks (GNNs). Some of them using supervised learning (Li et al., 2018; Sun & Yang, 2023; Li et al., 2024) and requires labelling training data using classical solvers. Alternatively, those based on reinforcement learning (Khalil et al., 2017; Ahn et al., 2020; Qiu et al., 2022; Sanokowski et al., 2023) and other unsupervised learning objectives (Karalias & Loukas, 2020; Sun et al., 2022; Zhang et al., 2023; Sanokowski et al., 2024) do not require labeled training data. Zhang et al. (2023) (LTFT) uses GFlowNets (Bengio et al., 2021a) which is related to reinforcement learning. Notably, Ahn et al. (2020); Sanokowski et al. (2023); Zhang et al. (2023) model MIS problem as a Markov Decision Process (MDP) and generate the solution step-by-step (autoregressively). While Sanokowski et al. (2023) fixes the order of node updates, Ahn et al. (2020) (LwD) and Zhang et al. (2023) (LTFT) choose which node to update at each step so that they have a “natural” or “pseudo-natural” serialization and most suitable for our analysis in Section 4.1 and Section C.3. In addition, Sun & Yang (2023) (DIFUSCO) and Sanokowski et al. (2024) (DiffUCO) both utilizes diffusion model.
Most of the algorithms above also work on other types of graph CO problems, such as Maximum Cut, Maximum Vetex Cover, and Minimum Dominating Set. Some (Khalil et al., 2017; Qiu et al., 2022; Sun & Yang, 2023) also work for the Travelling Salesman Problem (TSP).
Other GPU-based solvers for CO
Recently, some non-learning GPU-based solvers for CO problems have been developed. Sun et al. (2023) developed a GPU-accelerated sampling based method which works on MIS, Max Cut, and TSP. Schuetz et al. (2022); Ichikawa (2023) uses GNNs conduct non-convex optimization for MIS without machine learning. Alkhouri et al. (2024) uses direct quadratic optimization without GNNs for MIS.
Benchmarks for MIS
Böther et al. (2022) provides a benchmark for several MIS algorithms including Gurobi (Gurobi Optimization, LLC, 2024), KaMIS (ReduMIS(Lamm et al., 2017), Intel-Treesearch (Li et al., 2018), DGL-Treesearch, and LwD Ahn et al. (2020). It is the only MIS benchmark we know about including recent AI-inspired method, though it only focuses its comparsion for learning-based tree search algorithms. It suggests that LwD is better than learning-based tree search algorithms. This aligns with our results where LwD performs the best among learning-based algorithms. This benchmark covers various types of random graphs and several real-world datasets, so it is a good reference benchmark for comparison over different types of graphs. Though unlike our benchmark, it does not provide comparison across various size and density for random graphs. Our benchmark fills this gap and provides a more detailed comparison. We also include many newer AI-inspired algorithms, and greedy algorithms which leads to detailed analysis like serialization. Angelini & Ricci-Tersenghi (2023) provides a comparison between Schuetz et al. (2022) and Deg-Greedy over regular graphs. The algorithm papers also report experiments comparing with previous algorithms, but those comparisons usually only focus on a few datasets and a few selected baselines.
6 Conclusion and Takeaways
Given the great interest in designing “general purpose AI reasoners”, it is interesting to check how well recent AI-based methods have fared in combinatorial optimization, a field with a long history of ingenious hand-designed algorithms. Our careful empirical comparisons of such AI-inspired methods with classical methods on MIS problem showed that none of the new methods outperform ReduMIS, the best CPU-based MIS solver, which builds up the independent set iteratively, sometimes backtracking (i.e. delete a vertex from the current set). As the graphs get larger or denser, the superiority of ReduMIS becomes more evident, whereas several AI-inspired algorithms drop to performing no better than trivial classical algorithms such as Deg-Greedy.
Further analysis shows that the fact that AI-inspired algorithms like DIFUSCO and PCQO select the independent set in one shot may be handicapping themselves by foregoing the benefits of local search. Methods like Deg-Greedy and LTFT select one node at a step and do not backtrack end up performing weakly. Interestingly, the best-performing learning-based algorithm LwD is doing something in the middle: selecting several nodes at a time, which may allow a more effective approximation to local search.
Using local search as post-processing improves the solutions of AI-inspired methods more than the CPU-based algorithms such as degree-based greedy. Despite that, with the local search step, these AI-inspired methods still perform worse than ReduMIS. This suggests that a more promising method to use AI and machine learning techniques for combinatorial optimization might be to teach the model the value of local search instead of simply use it as a post-processing.
Acknowledgement
YW, HZ, and SA acknowledge funding from DARPA. The authors would like to thank Alvaro Velasquez, Ismail Alkhouri, Kaifeng Lyu, Sadhika Malladi, and Pravesh Kothari for helpful discussions.
Appendix

### v2

@@ 4 Performance Gap: Classical Methods Outperform AI-inspired Methods
Another key finding is that the simplest degree-based greedy (Deg-Greedy) serves as a remarkably strong baseline. As shown in Table 1, leveraging neural networks for node selection, LTFT often perform comparably to Deg-Greedy, particularly on larger or denser graphs. For example on ER graphs when $n=1000$ or $3000$, LTFT gives performance within 2% of Deg-Greedy (Table 1). Additionally, DIFUSCO and PCQO fail to outperform Deg-Greedy on larger or denser graphs, such as $n=1000$ with $d\geq 100$ and $n=3000$ for DIFUSCO, $n=3000$ with $d\geq 300$and $n=10000$ for PCQO.
Similar trends on other types of graphs
We also tested on Barabási–Albert (BA) graphs (Table 5), RB graphs (Table 6), and real-world graphs (Table 7). For BA graphs and RB graphs, ReduMIS also outperform all other methods across various setting. For real-world graphs, most methods including ReduMIS perform close to the best, likely because these datasets are either small or sparse and thus easy for MIS problem. On BA and RB graphs, we also observe that the performance gap between ReduMIS and AI-inspired methods increase as the graphs become larger and denser.
Deg-Greedy is also a strong baseline for all these graph types. Notably, for large RB graphs, all learning-based methods perform worse than Deg-Greedy. Since RB graphs are considered difficult cases for MIS (zhang2023let, ), these results may suggest learning-based methods are even weaker for graph types with higher "intrinsic hardness" than ER graphs.
5 Deconstructing the Shortfall: Comparison with Deg-Greedy
Our results show that the state-of-the-art AI-inspired algorithms for MIS still do not outperform the best heuristic ReduMIS. The surprising finding was that they also often do not outperform the simplest classical heuristic, Deg-Greedy, especially on large and dense graphs. In this section, we delve deeper into this comparison (Sections 5.1 and 5.2). We also explore the impact of augmenting various algorithms with a local search as post processing to enhance solution quality (Section 5.3).
@@ 5.1 Comparison between Deg-Greedy and LTFT
(a) Percentage of rounds when LTFT selects the node with smallest possible degree, i.e., behaves similarly to degree-based greedy. On ER graphs when the graph is dense (closer to the bottom left corner), the percentage of rounds LTFT selects the node with smallest possible degree is higher, i.e., behaves more similarly to degree-based greedy.
(b) The percentage to choose the smallest possible degree node on different part of the (degree-based) serialization. We find that the best algorithms (OnlineMIS, ReduMIS, iSCO) share a similar characteristic pattern, that they have high consistency with degree-based greedy on second and third part of the serialization, while on the first part there is a low consistency. PCQO and LwD, which performs a bit worse, also share this pattern but a bit higher for the first part, and lower for second and third parts. By contrast, LTFT and DIFUSCO, which perform significantly worse, either has very high consistency for all parts(LTFT) or has low consistency for all parts (DIFUSCO).
Figure 1: Analysis on ER graphs in (a) Section 5.1 and (b) Section 5.2
Deg-Greedy sequentially picks nodes for the independent set. At each step, it picks the node with the smallest degree in the residual graph (where the nodes in the independent set and their neighbors are removed). It does not reverse any decisions (ie once picked, the node stays in the independent set). As in Section 3.2 we call it a non-backtracking algorithm. LTFT is also a non-backtracking algorithm and it often perform similarly to Deg-Greedy in Tables 1 and 5. It uses a trained policy network GFlowNets (bengio2021flow, ) to pick a node for the independent set at each step. Thus, we can naturally compare LTFT with Deg-Greedy by investigating how close this trained policy compares to the naive policy in Deg-Greedy.
The results are shown in Figure 1(a). Overall, we observe that LTFT frequently selects nodes with very small degrees. On ER graphs where the average degree is at least 30, LTFT picks the node with the smallest degree over 85% of rounds except for sparse graphs ($d=10$ and $(n,d)=(3000,30)$). Moreover, in cases where LTFT selects the smallest degree nodes less frequently, its performance is worse than Deg-Greedy.
@@ 5.2 Serialization: allows comparing to Deg-Greedy
To evaluate the algorithms, we compute the percentage of rounds in which the smallest degree node is selected during serialization, similar to our comparison between LTFT and Deg-Greedy. Due to random tie-breaking in Algorithm 1, we repeat the process 1000 times and select the serialization with the highest number of rounds selecting the smallest degree node. In order to provide more fine-grained analysis, instead of reporting the overall percentage, we divide the serialization into three equal parts and report the percentages for each part.
Although Deg-Greedy theoretically achieves $100\%$ smallest-degree selections in its best serialization, there are exponential number of possible sequences and we cannot perfectly recover this with 1000 repetitions with random tie-breaking for all graphs. Nevertheless, for most graphs, Deg-Greedy has $100\%$ or very close to $100\%$ smallest-degree selections in our serialization.
The results, shown in Figure 1(b), compare the percentage of smallest-degree node selections across different algorithms. Due to space constraints, we present results only for ER graphs under selected parameters; additional results for ER graphs and BA graphs are available in Section C.4.
Algorithms that often perform the best, namely OnlineMIS, ReduMIS, and iSCO, exhibit a consistent pattern after serialization. In the first one-third of the serialization, these algorithms deviate significantly from selecting the smallest-degree nodes. However, in the middle and final thirds, the percentage of smallest-degree node selections increases substantially. This suggests that while degree-based greedy heuristics may appear shortsighted initially, they are highly effective in the later stages of solution construction. Interestingly, LwD and PCQO, which performs worse but close, also shares this pattern, but has higher percentage for the first part, but lower for second and third parts.
The similar pattern observed among several good-performing algorithms with very different techniques suggest intrinsic similarities among good solutions, regardless the algorithm leads to them. Analysis techniques targeting solutions, like serialization here, are especially suitable for discovering characteristics of good solutions, which can provide often overlooked insights for future algorithmic design.
By contrast, LTFT has very high percentage of smallest-degree nodes selection for all three parts. DIFUSCO has low percentage for all three parts. Both perform significantly worse than the best-performing algorithms. This suggests that these methods might fail to utilize (DIFUSCO) or emphasize too much (LTFT) on simple yet highly effective heuristics, such as greedily selecting the smallest-degree node, which may partly explain their performance limitations.
In Section C.5, we perform a pseudo-natural serialization for LwD. LwD is also a non-backtracking algorithm like LTFT, but it does not have a “natural serialization" like LTFT (Section 5.1) because it chooses several nodes in a step. The pseudo-natural serialization performs serialization in each step of LwD. The results in Figure 5 align with our “counterfactual" serialization results here.
@@ 5.3 Incorporating local search to improve solution
In the previous sections, we show that solutions generated by AI-based algorithms generally differ from those produced by degree-based greedy methods, which may explain their inferior performance on MIS problems. A natural idea to improve them would be to allow post-porcessing their solutions with local search. Some previous studies found this beneficial (ahn2020learning, ; boether_dltreesearch_2022, ).
@@
Table 2 shows the result of applying $2$-improvement local search (andrade2012fast, ), which is used in KaMIS, as a post-processing step for all algorithms (except OnlineMIS and ReduMIS since they already has local search). Full results for both ER and BA graphs are in Table 9 and details are in Section B.7.
Table 2: Incoperating local search as a post-processing procedure on selected ER graphs. We add $2$-improvement local search for the algorithm tested. We report the performance after the local search and the improvement from local search in (). Full results for ER and BA graphs in Table 9.
@@
Heuristics GPU-acc Learning-based $(n,d)$ Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO 1000,30 152.75 (1.75) 158.88 163.75 163.50 (0) 159.13 (0.5) 158.62 (0.24) 151.38 (1.38) 150.62 (6.87) 1000,100 60.75 (0.13) 64.75 66.63 66.50 (0) 60.88 (0.75) 64.12 (0.24) 61.88 (1.00) 57.88 (2.50) 3000,30 456.12 (4.24) 480.88 493.13 491.62 (0) 469.75 (5.5) 474.00 (0.75) 454.38 (5.38) 442.75 (29.37)
@@
As shown in Table 2, DIFUSCO benefit significantly more from the local search post-processing compared to others, such as Deg-Greedy and LwD. This observation aligns with our earlier findings: If the solution after serialization exhibits a high percentage of smallest-degree node selections in the later stages, there is relatively little room for improvement through local search. Conversely, if the solution after serialization shows a low percentage of smallest-degree node selections, there is greater potential for improvement via local search. In addition, although all algorithms except iSCO have improvements after local search, they still perform worse than ReduMIS.
6 Call for New Directions: Issues and Solutions
Building on our empirical findings and analysis on MIS as a testbed, we highlight several critical issues in the current research landscape of AI-inspired methods for combinatorial optimization (CO), and propose actionable directions to address them. We expect these issues and solutions are relevant for a wide range of CO problems, given similar challenges observed in tasks like TSP (joshi2021learning, ; sun2023difusco, ; xia2024position, ) and theoretical barriers identified for GNN-based approaches (gamarnik2023barriers, ). The following issue-solution pairs aim to guide future work toward more robust, insightful, and effective progress in this field.
Issue 1: Limited benchmarks and biased evaluation
Recent research on AI-inspired methods for combinatorial optimization (CO) often relies on limited test datasets and focuses on small-scale instances with tight inference time constraints, which can artificially favor AI-based methods (boether_dltreesearch_2022, ; ahn2020learning, ; sun2023difusco, ). While state-of-the-art classical solvers typically require more time, simple heuristics (like degree-based greedy for MIS), which are fast and possibly competitive on small instances, are often overlooked. Furthermore, current studies overlook that AI-inspired methods, especially on larger graphs, demand much higher computational resources—often exceeding 80GB of GPU memory for ER graphs with $10000$ nodes (Table 1), making them infeasible without multi-GPU setups. By contrast, KaMIS runs on a single CPU thread, scales efficiently up to millions of nodes (Table 8), and can be parallelized. Thus, classical solvers are often more scalable and practical than suggested by current benchmarks, which may not reflect true algorithmic strengths and weaknesses. Previous works (joshi2021learning, ; sun2023difusco, ; xia2024position, ) have highlighted problematic evaluations of neural-based methods on TSP, suggesting similar issues throughout CO.
Solution: Rigorous and comprehensive benchmarking
A critical step forward is to benchmark AI-inspired methods against classical solvers on a wider range of CO problems across varied instance sizes and realistic runtime settings. Comprehensive benchmarks, like (boether_dltreesearch_2022, ), as well as detailed experiments (Section 4) and analyses (Section 5) in our paper, are necessary for meaningful comparisons. Future evaluations should also report computation resources used (e.g., GPU memory) and allow for flexible inference times, ensuring a fair assessment across methods.
Issue 2: Intrinsic hardness and limits of learning
Our experiments (Section 4) show that learning-based methods consistently underperform leading classical solvers, even on in-distribution data thought to be ideal for ML. This suggests that problems like MIS may be fundamentally hard for current learning paradigms. Theoretical works highlight barriers for local and GNN-based algorithms, such as the Overlap Gap Property (OGP) (gamarnik2014limits, ; gamarnik2021overlap, ; gamarnik2023barriers, ).
Solution: Theoretical and empirical understanding of limitations
Addressing these limitations requires a combination of theoretical and empirical approaches. Future research should deepen understanding of theoretical barriers like OGP and develop empirical tools, such as serialization analysis (Section 5), to pinpoint failure modes of AI-inspired methods. This dual approach will clarify what is fundamentally difficult for learning in CO and guide new algorithmic developments.
Issue 3: Insufficient use and understanding of classical heuristics
Current AI-inspired algorithms for CO problems either attempt to fully replace classical heuristics or substitute classical components with learned models, often without understanding why classical methods succeed. Previous works (boether_dltreesearch_2022, ; xia2024position, ) reveal that the seeming success of some hybrid approaches, including neural network-guided tree search for MIS (li2018combinatorial, ) and TSP (xia2024position, ), typically derive most of their gains from embedded classical reductions rather than from the learned components.
Solution: Learning from and integrating classical heuristics
Progress demands that researchers move beyond treating classical solvers as black-box baselines. Instead, AI researchers tackling CO problems should analyze and learn the effective components of classical heuristics—such as reduction (xiao2013confining, ; akiba2016branch, ), local search (andrade2012fast, ), and evolutionary algorithms (lamm2015graph, ). It is also important to use empirical ablation studies to identify key ingredients, as done in (boether_dltreesearch_2022, ) and Section 5.3. Results in Section 5.2 show similarities among good solutions obtained from vastly differnt algorithms. Furthermore, our analysis Section 5 highlights that rather than relying solely on end-to-end methods like DIFUSCO, integrating classical heuristics such as greedily selecting the smallest-degree node into the overall algorithm may yield better results. One potential approach could involve using machine learning algorithms to identify a small subset of nodes, followed by a degree-based greedy method to complete the solution. These insights should then be systematically incorporated into AI-inspired frameworks, enabling principled hybrid algorithms that embed classical heuristics as core components.
7 Alternative Views
While our paper primarily targets algorithm designers and emphasizes the need to rethink and design better AI-inspired methods for combinatorial optimization problems, users who need to decide which method to use may hold different perspectives. Although our analysis clearly shows that AI-inspired methods currently lag behind classical heuristic solvers like KaMIS in terms of solution quality, there are situations where AI-inspired methods may be preferable. Learning-based AI methods often offer shorter inference times and can achieve comparable, or occasionally better, results when computation time is limited (boether_dltreesearch_2022, ; ahn2020learning, ; zhang2023let, ; alkhouri2024dataless, ). In scenarios where speed is prioritized over optimality, or where strict inference-time constraints exist, current AI-inspired approaches may be preferable options.
Nevertheless, it is still important to compare these methods against classical fast heuristics in time-sensitive settings. For example, in the MIS problem, the simple degree-based greedy algorithm also provides fast computation and often competitive performance, as shown in Section 4, and should be considered a strong baseline. Additionally, classical methods like KaMIS and degree-based greedy typically require fewer computational resources, such as GPU memory, than AI-inspired methods. Ultimately, there is no single “best” algorithm for the MIS problem—or for other combinatorial optimization problems—when practical use cases are considered. Users should select methods based on their specific requirements and constraints.
Appendices